### v1

@@ 4 Proposed Approach
In this paper, we study the case where $\mathcal{D}_{Z}$ is multivariate Gaussian. Specifically, we choose $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d})$, $i\in[1:n]$, with the all zeros vector of size $d\times 1$ denoted by $\mathbf{0}_{d}$ and the identity matrix of size $d\times d$ denoted by $\mathsf{I}_{d}$. The $k$th element of $\mathbf{Z}_{i}$ is denoted by $Z_{i,k}$ for $i\in[1:n]$ and $k\in[1:d]$. $Z_{i,k}$’s are distributed as, $[Z_{1,k},\dotsc,Z_{n,k}]^{T}\sim\mathcal{N}(\mathbf{0}_{n},\Sigma)$, for $k\in[1:d]$, where $\Sigma_{i,i}=\sigma^{2}$ and $\Sigma_{i,j}=r$ for $i,j\in[1:n]$, $i\neq j$, i.e.,
@@
$\begin{pmatrix}Z_{1,k}\\ Z_{2,k}\\ \vdots\\ Z_{n,k}\end{pmatrix}\sim\mathcal{N}\left(\mathbf{0}_{n},\begin{pmatrix}\sigma^{2}&r&\dotsc&r\\ r&\sigma^{2}&\dotsc&r\\ \vdots&\vdots&\vdots&\vdots\\ r&r&\dotsc&\sigma^{2}\end{pmatrix}\right),\quad k\in[1:d]$ (9)
The correlation coefficient is denoted by $\rho=\frac{r}{\sigma^{2}}$. Moreover, we let $\mathbb{E}[Z_{i,k}Z_{j,k^{\prime}}]=0$, $\forall i,j$, $\forall k\neq k^{\prime}$. We denote this class of distributions as $\mathcal{D}_{Z}^{G}$. Next, we present the noise generation protocol used to achieve the exact covariance structure among the privacy mechanisms of the distributed users, specified by $\Sigma$. Based on the values of $n,t,c,\epsilon,\delta$ considered in the system, the optimum noise parameters $\sigma^{2}_{*}$ and $r_{*}$ are calculated first (see Section 5 for optimum values). Then, the following steps generate the explicit noise terms with the required correlation structure.
  1. 1.
    Each pair of users $(i,j)$, $i,j\in[1:n]$, $i\neq j$ generates a pairwise random seed using the Diffie-Hellman key exchange.
  2. 2.
    Using the common seed, each pair of users $(i,j)$ samples the same random vector $\mathbf{S}_{i,j}\in\mathbb{R}^{d}$ from $\mathcal{N}(\mathbf{0}_{d},-r_{*}\mathsf{I}_{d})$,***We use use negative correlation among the privacy mechanisms of different users. Hence, $-r_{*}>0$. where $\mathbf{S}_{i,j}=\mathbf{S}_{j,i}$, and $\mathbf{S}_{i,j}$, $\mathbf{S}_{i^{\prime},j^{\prime}}$ are independent for any $(i,j)\neq(i^{\prime},j^{\prime})$.
  3. 3.
    Each user $i$ independently generates another noise variable $\mathbf{N}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}_{*}+r_{*}(n-1)\mathsf{I}_{d})$.
  4. 4.
    The combined noise added to $\mathbf{x}_{i}$ $i\in[1:n]$ is: $\mathbf{Z}_{i}=\sum_{j=1,j<i}^{n}\mathbf{S}_{i,j}-\sum_{j=1,j>i}^{n}\mathbf{S}_{i,j}+\mathbf{N}_{i}$. Then user $i$, $i\in[1:n]$ sends $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$ to the server.
5 Main Results
In this section, we provide the solution to (6) while satisfying (3), for the class of multivariate Gaussian distributions $\mathcal{D}_{Z}^{G}$ specified in Section 4. The following notation is used throughout the paper.
@@ 5 Main Results
$\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})=\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}_{2}\right]$ (10)
@@
Whenever $t=n$ (no dropouts) or $c=0$ (no collusion), we drop the $\mathcal{F}$ and $\mathcal{C}$ terms in the MSE notation in (10). In Proposition 1, we state the optimum decoder, $\bm{\alpha}^{*}$, (with $\alpha_{i}^{*}$ as its $i$th component) that minimizes the MSE for any given $\mathcal{F},\mathcal{C},\sigma^{2}$ and $\rho$.
@@
For any fixed $\mathcal{F},\mathcal{C}\subset[1:n]$ satisfying $|\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n$, and for any $\sigma^{2}$, $\rho$, the optimum decoder is given by,
$\bm{\alpha}^{*}=\arg\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},r,\bm{\alpha})=\frac{1}{1+\frac{d\sigma^{2}}{|\mathcal{U}|}(1+\rho(|\mathcal{F}|-1))}\mathbf{1}_{|\mathcal{U}|}$ (11)
Theorem 1 characterizes the MMSE in (6) for $\mathcal{D}_{Z}=\mathcal{D}_{Z}^{G}$ while ensuring the privacy constraint in Definition 1, for any $t$ and $c$. Note that finding the optimum $\mathcal{D}_{Z}^{G}$ in (6) is equivalent to optimizing $r$ (or equivalently $\rho$) and $\sigma^{2}$ that minimizes the MSE while satisfying the privacy constraint in Definition 1, based on the structure of $\mathcal{D}_{Z}^{G}$ specified in Section 4.
Theorem 1
For any given $\epsilon>0$, $\delta\in(0,1)$, $t$ and $c$, the optimum $\sigma^{2}$ and $\rho$ that solves (6) for $\mathcal{D}_{Z}=\mathcal{D}_{Z}^{G}$ while satisfying (3) is characterized by,
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})$
$\sigma^{2}_{*}$ =$=\begin{cases}\infty,&t=n\\ \frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}+\frac{\sigma^{2}_{\epsilon,\delta}(n-c-1)(n+c-2nc+t(n+c-2))}{(n-c)^{2}\sqrt{(t-c)(n-t)(n-c-1)}},&c<t<n.\end{cases}$ (12)
$\rho_{*}$ =$=\begin{cases}-\frac{\sigma^{2}_{*}-\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c}{2(n-1)(c-1)}+\frac{\sqrt{\left((n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c\right)^{2}+4(n-c-1)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)}}{2(n-1)(c-1)},&c\neq 1\end{cases}$ (13)
@@
$\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\geq|\mathcal{F}|\leq n\end{subarray}}$ $\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\mathcal{C},\sigma^{2},\rho,\bm{\alpha})=\left(1+\frac{(t-c)/d}{\sigma^{2}_{*}+\rho_{*}(t-1)}\right)^{-1}$ (14)
For a fixed $\epsilon$ and $\delta$, the value of $\sigma^{2}_{\epsilon,\delta}$ in Theorem 1 is the minimum variance of the Gaussian noise added to achieve $(\epsilon,\delta)$-DP in the standard Gaussian mechanism with an $L_{2}$ sensitivity of $2$. For any $\epsilon,\delta,t$ and $c$, Theorem 1 shows that $\rho_{*}\leq 0$ always holds. This implies that the (anti) correlated Gaussian mechanism outperforms (or performs equally when $c=t-1$) the independent Gaussian mechanism for any $\epsilon,\delta,t,c$. Next, we upper bound $\sigma^{2}_{\epsilon,\delta}$ using the results from [39, 40, 8] to better interpret the dependency of the MSE on $\epsilon$ and $\delta$. Generally, we assume that $\delta=10^{-5}$.
Proposition 2
@@
In Corollaries 13, we consider special cases of Theorem 1, and provide simplified MSE results using the bounds in Proposition 2.
@@
For $t=n$, $c=0$, the minimum MSE is given by,
$\min_{\sigma^{2},\rho,\bm{\alpha}}\mathrm{MSE}(n,\sigma^{2},\rho,\bm{\alpha})$ =$=\begin{cases}O\left(\frac{d}{n^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right),&\text{if }n^{2}>>d\\ O(1),&otherwise.\end{cases}$
@@
For $t<n$ and $c=0$, the minimum MSE is given by,
$\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{F}\subset[1:n]\\ t\leq|\mathcal{F}|<n\end{subarray}}$ $\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{F},\sigma^{2},\rho,\bm{\alpha})=\begin{cases}O\left(\frac{d(n-t)}{tn}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right),&\text{if }\frac{n^{2}}{n-t}>>d\\ O(1),&otherwise.\end{cases}$ (16)
@@
For $t=n$ and $c>0$ the minimum MSE is given by,
@@
$\min_{\sigma^{2},\rho}\quad\max_{\begin{subarray}{c}\mathcal{C}\subset[1:n]\\ 0\leq|\mathcal{C}|\leq c\end{subarray}}\min_{\bm{\alpha}}\mathrm{MSE}(n,\mathcal{C},\sigma^{2},\rho,\bm{\alpha})=\begin{cases}O\left(\frac{d}{(n-c)^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right),&\text{if }(n-c)^{2}>>d\\ O(1),&otherwise.\end{cases}$
@@
For the case of no dropouts (with or without colluding users) CorDP-DME achieves the same order of MSE as CDP (and DP-DME with SecAgg) for a fixed $\delta$, as shown in Corollaries 1 and 3. Note that for $|\mathcal{U}|$ non-colluding and responding users, CorDP-DME achieves a reduction in the MSE by a factor of $O(|\mathcal{U}|)$, compared to LDP. For the case of dropouts, CorDP-DME achieves MSEs that lie in between LDP and CDP (Corollary 2). If the maximum number of dropouts is $O(n^{p})$ for any $p<1$, Corollary 2 implies that the minimum MSE is $O(d\sigma^{2}_{\epsilon,\delta}/(tn^{1-p}))$, which has a scaling advantage over the case with independent noise, i.e., $\rho=0$, with a minimum MSE of $O(d\sigma^{2}_{\epsilon,\delta}/t)$.
@@
The optimum decoder in Proposition 1 results in a biased estimate of the mean. By fixing the decoder to be $\bm{\alpha}=\mathbf{1}_{|\mathcal{U}|}$, we can optimize the noise distribution and obtain an unbiased estimate with MSEs that lie in between the corresponding MSEs of DME with LDP and CDP. Proposition 3 characterizes the optimum noise distribution and the resulting MSE (unbiased) for the case where $\bm{\alpha}=\mathbf{1}_{|\mathcal{U}|}$.
Proposition 3
For any $\epsilon>0$ and $\delta\in(0,1)$, the $r_{*}$ and $\sigma^{2}_{*}$ given in (13) and (12) satisfies,
@@
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ |\mathcal{C}|\leq c<t\leq|\mathcal{F}|\leq n\end{subarray}}\sup_{\mathbf{x}_{j}\in\mathbb{B}^{d},j\in\mathcal{U}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}_{2}\right],$ (17)
and the resulting MSE is given by $\frac{d(\sigma^{2}_{*}+r_{*}(t-1))}{t-c}$.

### v2

@@ IV Proposed Approach: CorDP-DME Protocol
In this paper, we study the case where $\mathcal{D}_{Z}$ in (7) is multivariate Gaussian. Specifically, we choose,
$\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}\mathsf{I}_{d}),\quad i\in[1:n]$ (14)
with the all zeros vector of size $d\times 1$ denoted by $\mathbf{0}_{d}$ and the identity matrix of size $d\times d$ denoted by $\mathsf{I}_{d}$. The $k$th element of $\mathbf{Z}_{i}$ is denoted by $Z_{i,k}$ for $i\in[1:n]$ and $k\in[1:d]$. The $Z_{i,k}$s are distributed as:
@@
$\begin{pmatrix}Z_{1,k}\\ Z_{2,k}\\ \vdots\\ Z_{n,k}\end{pmatrix}\sim\mathcal{N}\left(\mathbf{0}_{n},\sigma^{2}\begin{pmatrix}1&\rho&\dotsc&\rho\\ \rho&1&\dotsc&\rho\\ \vdots&\vdots&\vdots&\vdots\\ \rho&\rho&\dotsc&1\end{pmatrix}_{n\times n}\right)$ (15)
[…]
Next, we provide the main results of this paper, that characterize the optimal encoding decoding functions and the corresponding minimum MSEs.
IV-A Main Results
In this section, we provide the solution to (7) while satisfying the privacy constraint in Definition 1, for the class of multivariate Gaussian distributions $\mathcal{D}_{Z}^{G}$ specified in (15). We present the optimum decoder and the optimum noise distribution, along with the corresponding minimum MSE for the general case with arbitrary $n,t,c,\epsilon,\delta$ in Theorem 1. Subsequently, in Corollaries 1-3, we derive the MSEs of the special cases, with specific $t$ and $c$. Furthermore, Theorems 2 and 3 provide further achievability and converse results for unbiased estimates. The following notation is used throughout the paper.
$\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$
@@ IV-A Main Results
$=\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\alpha_{i}(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$ (16)
@@
In Proposition 1, we state the optimum decoder, $\boldsymbol{\alpha}^{*}_{\mathcal{U}}$, (with $\alpha_{i}^{*}$ as its $i$th component) that minimizes the MSE for any set of responding users $\mathcal{U}$, and any given noise distribution characterized by $\sigma^{2}$ and $\rho$.
@@
For any fixed $\mathcal{U}\subseteq\mathcal{U}_{all}$ satisfying $|\mathcal{U}|\geq t$, and for any $\sigma^{2}$, $\rho$, the optimum decoder is:
$\boldsymbol{\alpha}^{*}_{\mathcal{U}}$ $=\arg\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$ (17)
$=\frac{1}{1+\frac{d\sigma^{2}}{|\mathcal{U}|}(1+\rho(|\mathcal{U}|-1))}\mathbf{1}_{|\mathcal{U}|}.$ (18)
The corresponding MSE is given by,
$\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})\!=\!\left(1\!+\!\frac{|\mathcal{U}|/d}{\sigma^{2}(1+\rho(|\mathcal{U}|\!-\!1))}\right)^{-1}$ (19)
The proof of Proposition 1 is given in Appendix B. Next, we provide Theorem 1, which characterizes the MMSE in (7) for $\mathcal{D}_{Z}=\mathcal{D}_{Z}^{G}$ while ensuring the privacy constraint in Definition 1, for any $t$ and $c$. Note that finding the optimum $\mathcal{D}_{Z}^{G}$ in (7) is equivalent to optimizing $\rho$ and $\sigma^{2}$ that minimizes the MSE while satisfying the privacy constraint in Definition 1, based on the structure of $\mathcal{D}_{Z}^{G}$ specified in (15).
Theorem 1 (Optimum noise distribution)
For any given $\epsilon>0$, $\delta\in(0,1)$, $t\leq n$ and $c<t$, the optimum $\sigma^{2}$ and $\rho$ that solves (7) for $\mathcal{D}_{Z}=\mathcal{D}_{Z}^{G}$ while satisfying the generalized privacy constraint in Definition 1 for any $\mathcal{U}_{col}\subset\mathcal{U}_{all}$ with $|\mathcal{U}_{col}|\leq c$ is characterized by,
$(\sigma^{2}_{*},\rho_{*})=\arg\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$
$\sigma^{2}_{*}$ =$\!=\!\begin{cases}\infty,&t=n\\ \frac{\sigma^{2}_{\epsilon,\delta}(n^{2}-2n-cn+2)}{(n-c)^{2}}\\ +\frac{\sigma^{2}_{\epsilon,\delta}(n-c-1)(n+c-2nc+t(n+c-2))}{(n-c)^{2}\sqrt{(t-c)(n-t)(n-c-1)}},&c\!<\!t\!<\!n.\end{cases}$ (20)
$\rho_{*}$ =$\!=\!\begin{cases}-\frac{\sigma^{2}_{*}-\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}(n-1)-\sigma^{2}_{\epsilon,\delta}(n-2)},&c=1\\ \frac{-(n-2)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c}{2(n-1)(c-1)}\\ \quad+\frac{\!\sqrt{\left(\!(n-2)\left(\!1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\right)-c\!\right)^{2}\!\!+4(n-c-1)\left(1-\frac{\sigma^{2}_{\epsilon,\delta}}{\sigma^{2}_{*}}\!\right)}}{2(n-1)(c-1)},&c\neq 1\end{cases}$ (21)
@@
$\min_{\sigma^{2},\rho}\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}$ $\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$
$=\left(1+\frac{t/d}{\sigma^{2}_{*}(1+\rho_{*}(t-1))}\right)^{-1}$ (22)
The proof of Theorem 1 is given in Appendix C. For a fixed $\epsilon$ and $\delta$, the value of $\sigma^{2}_{\epsilon,\delta}$ in Theorem 1 is the minimum variance of the Gaussian noise added to achieve $(\epsilon,\delta)$-DP in the standard Gaussian mechanism with an $L_{2}$ sensitivity of $2$. For any $\epsilon,\delta,t$ and $c$, Theorem 1 shows that $\rho_{*}\leq 0$ always holds. This implies that the (anti) correlated Gaussian mechanism outperforms (or performs equally when $c=t-1$) the independent Gaussian mechanism for any $\epsilon,\delta,t,c$. Next, we upper bound $\sigma^{2}_{\epsilon,\delta}$ using the results from [45, 46, 8] to better interpret the dependency of the MSE on $\epsilon$ and $\delta$. Generally, we assume that $\delta=10^{-5}$.
Proposition 2 (Bounds on $\sigma^{2}_{\epsilon,\delta}$)
@@
The proof of Proposition 2 is given in Appendix D. In Corollaries 13, we consider special cases of Theorem 1, and provide simplified MSE results using the bounds in Proposition 2.
@@
For $t=n$, $c=0$, the minimum MSE while satisfying $(\epsilon,\delta)$-DP in Definition 1 with $\mathcal{U}_{col}=\emptyset$ is given by,
$\min_{\sigma^{2},\rho,\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}$ $(\sigma^{2},\rho,\mathcal{U}_{all},\boldsymbol{\alpha}_{\mathcal{U}_{all}})$
=$=\begin{cases}O\left(\frac{d}{n^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right),&\text{if }n^{2}>>d\\ O(1),&otherwise.\end{cases}$
@@
For $t<n$ and $c=0$, the minimum MSE while satisfying $(\epsilon,\delta)$-DP in Definition 1 with $\mathcal{U}_{col}=\emptyset$ is given by,
$\min_{\sigma^{2},\rho}$ $\max_{\begin{subarray}{c}\mathcal{U}\subset[1:n]\\ t\leq|\mathcal{U}|<n\end{subarray}}\min_{\boldsymbol{\alpha}_{\mathcal{U}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U},\boldsymbol{\alpha}_{\mathcal{U}})$
=$\quad=\begin{cases}O\left(\frac{d(n-t)}{tn}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right),&\text{if }\frac{nt}{n-t}>>d\\ O(1),&otherwise.\end{cases}$ (24)
@@
For $t=n$ and $c>0$ the minimum MSE while satisfying $(\epsilon,\delta)$-DP in Definition 1 for any $\mathcal{U}_{col}\subset\mathcal{U}_{all}$ with $|\mathcal{U}_{col}|\leq c$ is given by,
$\min_{\sigma^{2},\rho,\boldsymbol{\alpha}_{\mathcal{U}_{all}}}\mathrm{MSE}(\sigma^{2},\rho,\mathcal{U}_{all},\boldsymbol{\alpha}_{\mathcal{U}_{all}})$
@@
=$\quad=\begin{cases}O\left(\frac{d}{n(n-c)}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right),&\text{if }n(n-c)>>d\\ O(1),&otherwise.\end{cases}$ (25)
For the case of no dropouts and no collusions, CorDP-DME achieves the same order of MSE as CDP (and DP-DME with SecAgg) for any $\epsilon$ and $\delta$, as shown in Corollary 1. In the absence of dropouts, CorDP-DME reduces the MSE compared to LDP by a factor of $O(n)$ without collusion, and by $O(n-c)$ with collusion, as shown in Corollaries 1 and  3.
@@
For the case of dropouts, CorDP-DME achieves MSEs that lie in between LDP and CDP (Corollary 2 and Section V). If the maximum number of dropouts is $O(n^{p})$ for any $p<1$, Corollary 2 implies that the minimum MSE is $O\left(\frac{d\sigma^{2}_{\epsilon,\delta}}{tn^{1-p}}\right)$, which has a scaling advantage over the case with independent noise, i.e., $\rho=0$, that has a minimum MSE of $O\left(\frac{d\sigma^{2}_{\epsilon,\delta}}{t}\right)$.
@@
The optimum decoder in Proposition 1 results in a biased estimate of the mean. As we assume linear decoders $\boldsymbol{\alpha}_{\mathcal{U}}$ that are independent of $M(\mathbf{x}_{i})$, $\forall i$ and zero mean Gaussian noise for the privacy mechanism, an unbiased estimate is obtained if and only if $\boldsymbol{\alpha}_{\mathcal{U}}=\mathbf{1}_{\mathcal{U}}$ for any $\mathcal{U}\subseteq\mathcal{U}_{all}$, in (2). Theorem 2 characterizes the optimum noise distribution and the resulting MSE for the case of unbiased estimates.
Theorem 2 (Unbiased mean estimate)
Let $\tilde{\mathbf{S}}_{\mathcal{U}}=\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}(\mathbf{x}+\mathbf{Z}_{i})$ be an unbiased estimate of the true mean $\mathbf{S}_{\mathcal{U}}=\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}$, where $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ follows distributions from $\mathcal{D}_{Z}^{G}$. For any given $\epsilon>0$, $\delta\in(0,1)$, $t\leq n$ and $c<t$, the $\sigma^{2}_{*}$ and $\rho_{*}$ given in (20) and (21) satisfy,
@@
$(\sigma^{2}_{*},\rho_{*})$ $=\arg\min_{\sigma^{2},\rho}\quad\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}\quad\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}$
$\quad\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$ (26)
while satisfying $(\epsilon,\delta)$-DP in Definition 1 for any $\mathcal{U}_{col}\subset\mathcal{U}_{all}$ with $|\mathcal{U}_{col}|\leq c$. The resulting MSE is given by:
$\min_{\sigma^{2},\rho}\quad\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}$ $\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}$
$\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}(\mathbf{x}_{i}+\mathbf{Z}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$
$=\frac{d\sigma^{2}_{*}(1+\rho_{*}(t-1))}{t}$ (27)